OpenStudy (anonymous):

Explain the difference between using the tangent ratio to solve for a missing angle in a right triangle versus using the cotangent ratio. You must use complete sentences and any evidence needed (such as an example) to prove your point of view. (10 points)

OpenStudy (dumbcow):

hmm well the only difference is that cot is the reciprocal of tangent if tan = 3/5 then cot = 5/3

OpenStudy (dumbcow):

or cot = adj/opp

OpenStudy (anonymous):

so i'd just write ; Tan 3/5 than cot = 5/3 or cot = adj/opp ?

OpenStudy (dumbcow):

no please dont quote me lol

OpenStudy (anonymous):

i found 1 last question .

OpenStudy (anonymous):

Use the image below to answer the following question. Find the value of sin x° and cos y°. What relationship do the ratios of sin x° and cos y° share? (10 points)

OpenStudy (dumbcow):

"solve for missing angle" you have to take inverse tangent to get angle if you were to use cot then you would have to take reciprocal first, then take inverse tangent so difference is you would have 1 extra step if you used the cotangent ratio please put in your own words and come up with your own example

OpenStudy (dumbcow):

notice that x+y = 90 notice that sinx = cosy and siny = cosx

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

can you tell me what the value of sin x is and cos y ?

OpenStudy (dumbcow):

use SOH CAH TOA with triangle provided

OpenStudy (anonymous):

can you be more specific

OpenStudy (jdoe0001):

https://www.youtube.com/watch?v=qDH3qRz5wZE <---- maybe?

OpenStudy (anonymous):

can one of you just give me the value of sin x and cos y ?

OpenStudy (anonymous):

please !?

OpenStudy (dumbcow):

watch the link posted

OpenStudy (anonymous):

that didnt help . im just not gonna do this problem lol

OpenStudy (dumbcow):

didnt help? you havent had time to watch it and try it yet??

OpenStudy (anonymous):

i watched the video when he first sent it . it didn't help me solve my problem

OpenStudy (dumbcow):

it explains SOH CAH TOA and how you can find sin x and cos y on your own

OpenStudy (anonymous):

its okay . thanks tho

OpenStudy (jdoe0001):

well... draw a right triangle here.... give it some lengths and let's find the angle using tangent